# Predicting perturbation targets with causal differential networks

### v1

@@ 4.2 Synthetic experiments
There are two considerations that motivate further experiments in controlled settings. First, while it would be ideal to evaluate all algorithms on real data, current causal discovery algorithms that support unknown interventions are not tractable on datasets with more than tens of variables. In fact, many baselines require hours on even $N=10$ datasets, and they do not scale favorably (Table 9). Our transcriptomics datasets contain hundreds of genes, even after filtering to those that are differentially expressed (Figures 6 and 7). On the other hand, existing models for biological perturbations all rely on some form of domain knowledge, and their performance is inseparable from the choice and quality of these external data. Here, synthetic settings allow us to assess the model’s capacity to predict intervention targets in isolation.
Datasets
We generated 120 synthetic test datasets with hard and soft interventions. To generate observational data, we sampled Erdős-Rényi graphs with $N=10,20$ nodes and $E=N,2N$ expected edges; causal mechanism parameters; and observations of each variable, in topological order. We sampled $3N$ distinct subsets of 1-3 nodes ($N$ each) as intervention targets. For hard interventions, we set $x\leftarrow z$, where $z$ is uniform. To emulate transcriptomics data, in which perturbation effects are measured in fold change, we introduce soft interventions: $x\leftarrow cf(\pi_{x})$, where $c$ is a positive scaling factor, with equal probability $c\lessgtr 1$. For our synthetic training set, we generated approximately 4000 datasets following the training set of Sea, with hard interventions only. Full details are available in Appendix B.2.
Baselines
We compare against discrete and continuous causal discovery algorithms for unknown interventions. Ut-Igsp (Squires et al., 2020) infers causal graphs and unknown targets by greedily selecting the permutation of variables that minimizes their proposed score function. Dcdi (Brouillard et al., 2020) and Bacadi (Hägele et al., 2023) are continuous causal discovery algorithms that fit generative models to the data, where the causal graph and intervention targets are model parameters. The -G and -Dsf suffixes on Dcdi correspond to Gaussian and deep sigmoidal flow parametrizations of the likelihood. Bacadi is evaluated here using the fully-connected implementation (better performance), with the linear version in Table 11. Its -E and -M suffixes indicate empirical (standard) and mixture (bootstrap) variants.
We also compare against the difference causal inference (Dci) algorithm with stability selection (Belyaeva et al., 2021), a discrete optimization algorithm that aims to detect edge-level differences between two causal graphs. On this node-centric task, we take each node’s proportion of changed edges as its likelihood of being an intervention target. While Dci was also motivated by biological applications, it only scales to around a hundred nodes at most, so we evaluate it alongside other causal discovery methods here.
Evaluation
In the synthetic case, we are not constrained by biological redundancy or incomplete predictions, so we report standard classification metrics.
  • Mean average precision (mAP) is the average precision over a continuum of binarization thresholds, computed independently for each variable, averaged over all regimes of the same number of targets. This ranges from 0 to 1 (perfect), where the random baseline depends on the positive rate.
  • Area under the receiver operating curve (AUC) is also computed independently for each variable, averaged over the same set. This ranges from 0 to 1 (perfect), where the random baseline is 0.5 (per edge).
Results
Table 4: Intervention target prediction results on synthetic datasets with $N=10,E=10$. Uncertainty is standard deviation over 5 distinct datasets. Top: hard interventions; bottom: soft interventions (“scale”). Number in parentheses indicates number of intervention targets. Runtimes are documented in Table 9, and full results are available in Table 11.
@@ Results
Hard Ut-Igsp $0.19\scriptstyle\pm.03$ $0.56\scriptstyle\pm.05$ $0.38\scriptstyle\pm.04$ $0.57\scriptstyle\pm.05$ $0.27\scriptstyle\pm.05$ $0.63\scriptstyle\pm.03$ $0.38\scriptstyle\pm.03$ 0.55±.03plus-or-minus0.55.030.$5$03
Dcdi-G $0.38\scriptstyle\pm.12$ $0.59\scriptstyle\pm.11$ $0.47\scriptstyle\pm.03$ $0.53\scriptstyle\pm.03$ $0.27\scriptstyle\pm.09$ $0.47\scriptstyle\pm.07$ $0.45\scriptstyle\pm.04$ $0.49\scriptstyle\pm.05$
Dcdi-Dsf $0.32\scriptstyle\pm.05$ $0.58\scriptstyle\pm.10$ $0.43\scriptstyle\pm.03$ $0.48\scriptstyle\pm.03$ $0.25\scriptstyle\pm.04$ $0.42\scriptstyle\pm.05$ $0.46\scriptstyle\pm.03$ $50$05
Dci $0.42\scriptstyle\pm.07$ $0.75\scriptstyle\pm.03$ $0.63\scriptstyle\pm.11$ $80$08 $0.39\scriptstyle\pm.10$ $0.72\scriptstyle\pm.06$ $0.53\scriptstyle\pm.09$ $0.69\scriptstyle\pm.07$
BaCaDI-E $0.16\scriptstyle\pm.04$ $0.61\scriptstyle\pm.09$ $0.36\scriptstyle\pm.04$ $60$06 $0.13\scriptstyle\pm.02$ $0.62\scriptstyle\pm.05$ $0.36\scriptstyle\pm.03$ $0.60\scriptstyle\pm.04$
BaCaDI-M $0.12\scriptstyle\pm.01$ $0.57\scriptstyle\pm.05$ 0.33±.02plus-or-minus0.33.020.$3$02 $0.56\scriptstyle\pm.04$ $0.12\scriptstyle\pm.01$ $0.57\scriptstyle\pm.04$ 0.33±.02plus-or-minus0.33.020.$3$02 0.55±.04plus-or-minus0.55.040.$5$04
@@
Soft Ut-Igsp $20$02 $0.71\scriptstyle\pm.04$ $0.37\scriptstyle\pm.01$ $0.54\scriptstyle\pm.03$ $0.24\scriptstyle\pm.03$ $0.72\scriptstyle\pm.03$ $0.39\scriptstyle\pm.04$ $0.58\scriptstyle\pm.05$
Dcdi-G $0.38\scriptstyle\pm.11$ $0.58\scriptstyle\pm.11$ $0.48\scriptstyle\pm.05$ $0.50\scriptstyle\pm.07$ $0.25\scriptstyle\pm.06$ $0.49\scriptstyle\pm.08$ $0.48\scriptstyle\pm.05$ $0.53\scriptstyle\pm.06$
Dcdi-Dsf $0.37\scriptstyle\pm.09$ $0.59\scriptstyle\pm.11$ $0.45\scriptstyle\pm.05$ $0.51\scriptstyle\pm.06$ $0.32\scriptstyle\pm.09$ $0.53\scriptstyle\pm.13$ $0.43\scriptstyle\pm.04$ $0.47\scriptstyle\pm.04$
Dci $0.45\scriptstyle\pm.10$ $\textbf{0.73}\scriptstyle\pm.07$ $0.54\scriptstyle\pm.04$ $\textbf{0.68}\scriptstyle\pm.03$ $0.34\scriptstyle\pm.03$ $0.69\scriptstyle\pm.03$ $0.45\scriptstyle\pm.03$ $0.62\scriptstyle\pm.01$
BaCaDI-E $0.34\scriptstyle\pm.16$ $0.63\scriptstyle\pm.10$ $0.49\scriptstyle\pm.14$ $0.62\scriptstyle\pm.15$ 0.22±.03plus-or-minus0.22.030.$2$03 $0.64\scriptstyle\pm.05$ $0.38\scriptstyle\pm.06$ $0.58\scriptstyle\pm.06$
BaCaDI-M $0.28\scriptstyle\pm.11$ $0.62\scriptstyle\pm.10$ 0.44±.14plus-or-minus0.44.140.$4$14 $0.59\scriptstyle\pm.15$ $0.21\scriptstyle\pm.02$ $0.64\scriptstyle\pm.05$ $0.38\scriptstyle\pm.06$ $0.58\scriptstyle\pm.06$
Cdn (Mlp) $0.29\scriptstyle\pm.04$ $0.53\scriptstyle\pm.05$ 0.44±.04plus-or-minus0.44.040.$4$04 $0.47\scriptstyle\pm.02$ 0.33±.09plus-or-minus0.33.090.$3$09 $0.50\scriptstyle\pm.09$ $0.52\scriptstyle\pm.06$ $0.56\scriptstyle\pm.05$
[…]
On synthetic data, Cdn achieves high performance across intervention types and data-generating mechanisms (Table 4), while running in seconds (Table 9). As an ablation study, we investigated replacing the differential network with a multi-layer perceptron, which does not model interactions between edges. On the hard interventions, the Mlp is sufficient for predicting intervention targets, perhaps by abusing marginal variances (Reisach et al., 2021) or other node-level artifacts of synthetic data. Once we consider soft interventions, however, the original Cdn formulation significantly outperforms the Mlp variant – indicating that graph-level information is essential.
@@
Dci is the best baseline, which is encouraging, as it was designed for the edge-centric version of our task, with the intuition that detecting edge differences is easier than reproducing the entire causal graph. It performs particularly well on linear data, which align with its assumptions, and less well on non-linear data. Surprisingly, most other baselines perform poorly at recovering intervention targets. In the case of Dcdi and BaCaDI, this may be because it is hard to select a single sparsity threshold for varying sizes of intervention sets, and to balance the regularizer with the generative modeling objective.
5 Conclusion
We presented Cdn, a causal discovery-powered approach for predicting perturbation targets in the context of single-cell transcriptomics. Cdn uses an amortized causal discovery algorithm to represent a pair of observational and interventional datasets in terms of their data-generating mechanisms, and then trains a classifier to identify variables whose conditional independencies have changed. Our approach achieves the state-of-the-art in predicting perturbation targets over seven transcriptomics datasets, compared to a variety of perturbation modeling baselines. Cdn also surpasses causal discovery algorithms for predicting unknown targets across synthetic settings, while only requiring a fraction of the runtime. Nonetheless, there is substantial space for improvement, so we hope that this work and the associated datasets will inform future method development for these biological applications.
Reproducibility statement
All source code will be made publicly available. Hyperparameter and implementation choices are described in Section 3.3. Details regarding biological data pre-processing can be found in Section 4.1 and Appendix B.1. The raw data are freely available via the accession codes listed in Table 5. Synthetic data generation is discussed in Appendix B.2. Details required to reproduce baselines are described in Sections 4.1 and 4.2, as well as Appendix C.1.
Ethics statement
We use publicly available datasets which do not involve human subjects. While our work has applications to biological understanding, our work does not directly concern the design of molecules or other potentially harmful agents.
@@ Acknowledgements
Appendix A Theoretical intuition
While this paper focuses on biological applications, we would like to show that our model is well-specified, i.e. it has the capacity to predict correct intervention targets, given certain inputs.
@@ Appendix A Theoretical intuition
Continuous causal discovery works typically rely on the universality of their architectures for consistency (Brouillard et al., 2020), so this analysis also focuses on computational capacity. Theoretically, it is impossible to “guarantee” that high-dimensional representations from learned model contain or do not contain certain information. This can be tested empirically through probing, e.g. querying for grammatical structure in language models (Vulić et al., 2020), but cannot be “proven” in the traditional sense. Thus, we believe that the empirical performance of our model on real datasets is its primary contribution, but provide these explanations for interested readers. In this analysis, we assume that the “causal” representations $h$ contains enough information both to retain global statistics and recover the true graph $E$. The former is reasonable due to high model capacity, and the latter is based on high empirical performance in graph reconstruction (Wu et al., 2024). We emphasize that the latter is an empirical judgment, which may not hold on all datasets in practice.
Axial attention architecture
Our differential network is implemented using an axial attention layer, which is composed of two self-attention layers (one along each axis of the adjacency matrix) and a feed-forward network. For simplicity, we follow prior work Yun et al. (2019) and ignore layer normalization and dropout.
Our inputs $h\in\mathbb{R}^{2d\times N\times N}$ are $2d$-dimension features, which represent a pair of $N\times N$ causal graphs. We use $h_{\cdot,j}$ to denote a length $N$ row for a fixed column $j$, and $h_{i,\cdot}$ to denote a length $N$ column for a fixed row $i$. The axial attention layer implements:
@@
Here, the axial attention layer should implement $h-h^{\prime}$, so that when we collapse over the incoming edges, the output is non-zero only at the edge differences. Suppose the first dimension of the $2d$ feature stores $E$, and the second dimension stores $E^{\prime}$. The row self-attention implements the identity (in the first two dimensions). Then we can set $W_{K,Q}$ to zero, $W_{V}$ to the identity, and $W_{O}$ to
@@
where $c_{i}$ are sampled at random per synthetic dataset. This choice is inspired by the fact that biological perturbation effects are measured in fold-change. Here, the adjacency matrices are the same, but global statistics differ. In particular, we focus on two statistics: the correlation matrix $R$ and the covariance matrix $\Sigma$. Note that while these inputs differ slightly from our main experiments (reasons discussed in Appendix D), we show that they still achieve reasonable performance in Table 8.
@@
Finally, we re-iterate that Cdn operates over high-dimensional features, rather than the inputs and outputs themselves. Thus, these sketches only serve to provide context regarding the axial attention framework’s computational capacity, rather than as a blueprint for the computations it performs.
Appendix B Datasets
B.1 Biological datasets
We converted all single cell datasets to LogTP10K + 1 expression values (log-normalized, transcripts per 10,000 UMIs). Perturb-seq dataset variables represented genes that were mapped and filtered by the authors. Sci-Plex dataset variables represented genes that appeared in at least 5,000 cells (threshold chosen to achieve a similar number of genes). We performed differential expression analysis via the scanpy package (Wolf et al., 2018), using the Wilcoxon signed-rank test (Wilcoxon, 1945) with Benjamini-Hochberg p-value correction and a threshold of adjusted p-value $<0.05$. Table 5 reports statistics of the raw, unprocessed datasets.
For Perturb-seq datasets, we kept perturbations with $>10$ differentially-expressed genes (DEGs), and clustered them using k-means with $k=200$, chosen heuristically based on log-fold change heatmaps (Figure 5). These clusters were used to inform data splits for seen cell lines, where the largest cluster was allocated to the training set, and all remaining clusters were split equally among train and test. The largest cluster(s) appear to contain perturbations with smaller effects. For Sci-Plex datasets, only 6 drug perturbations across 2 cell lines resulted in differential expression of their known protein targets (Supplementary Table 8 from McFaline-Figueroa et al. (2024)). Therefore, we used these exclusively as test sets. Table 6 reports statistics of the final, processed datasets. Figures 6 and 7 plot the full distribution of number of cells and DEGs per perturbation.
The full names of the Sci-Plex drugs are as follows.
  • Infigratinib (“infig”)
  • Nintedanib (“nint”)
  • Palbociclib (“palb”)
  • Doxorubicin (“doxo”)
  • Volasertib (“vola”)

### v2

@@ 4.2 Synthetic experiments
While it would be ideal to evaluate all algorithms on real data, current causal discovery algorithms that support unknown interventions are not tractable on datasets with more than tens of variables. In fact, many baselines require hours on even $N=10$ datasets, and they do not scale favorably (Table 11). Our transcriptomics datasets contain hundreds of genes, even after filtering to those that are differentially expressed (Figure 8). At the same time, existing models for biological perturbations all rely on some form of domain knowledge, and their performance is inseparable from the choice and quality of these external data. Synthetic data allow us to assess the model’s capacity to predict intervention targets in isolation.
Datasets  Each evaluation setting consists of 5 unseen, i.i.d. sampled graphs. To assess the generalization capacity of our framework, we tested the model on seen (linear) and unseen causal mechanisms (polynomial, sigmoid), and unseen soft interventions. To generate observational data, we sampled Erdős-Rényi graphs with $N=10,20$ nodes and $E=N,2N$ expected edges; causal mechanism parameters; and observations of each variable, in topological order. We sampled $3N$ distinct subsets of 1-3 nodes ($N$ each) as intervention targets. Hard interventions set $x\leftarrow z$, where $z$ is uniform. For soft interventions (Section 3.3), we trained on shift and tested on scale for synthetic experiments (see Table 9 for explanation). Real experiments trained on all interventions.
Table 3: Intervention target prediction results on synthetic datasets with $N=20,E=40$. Top: hard interventions (uniform); bottom: soft interventions (scale). Neither polynomial nor “scale” soft interventions were seen during training. Number in parentheses indicates number of intervention targets. Uncertainty is standard deviation over 5 i.i.d. datasets. Runtimes in Table 11. Extended results in Table 10.
@@
Hard Ut-Igsp $10$01 $0.63\scriptstyle\pm.03$ $0.18\scriptstyle\pm.01$ $0.52\scriptstyle\pm.03$ $0.19\scriptstyle\pm.02$ $0.70\scriptstyle\pm.02$ $0.23\scriptstyle\pm.03$ $0.56\scriptstyle\pm.03$
Dcdi-G $0.16\scriptstyle\pm.02$ $0.50\scriptstyle\pm.02$ $0.27\scriptstyle\pm.05$ $0.50\scriptstyle\pm.04$ $0.15\scriptstyle\pm.03$ $0.45\scriptstyle\pm.04$ $0.26\scriptstyle\pm.02$ $0.51\scriptstyle\pm.04$
Dcdi-Dsf $0.15\scriptstyle\pm.02$ $0.46\scriptstyle\pm.03$ $0.26\scriptstyle\pm.03$ $0.47\scriptstyle\pm.04$ $0.18\scriptstyle\pm.05$ $0.48\scriptstyle\pm.08$ $0.28\scriptstyle\pm.04$ $0.52\scriptstyle\pm.05$
Bacadi-E $0.18\scriptstyle\pm.09$ $0.71\scriptstyle\pm.06$ $0.28\scriptstyle\pm.02$ $0.69\scriptstyle\pm.03$ $0.20\scriptstyle\pm.06$ $0.84\scriptstyle\pm.04$ $0.37\scriptstyle\pm.07$ $0.82\scriptstyle\pm.05$
Bacadi-M $0.10\scriptstyle\pm.02$ 0.66±.05plus-or-minus0.66.050.$6$05 $0.21\scriptstyle\pm.01$ $0.63\scriptstyle\pm.02$ $0.09\scriptstyle\pm.02$ $0.71\scriptstyle\pm.05$ $0.24\scriptstyle\pm.03$ $0.70\scriptstyle\pm.05$
Dci $0.58\scriptstyle\pm.07$ $0.85\scriptstyle\pm.04$ $0.63\scriptstyle\pm.06$ $0.84\scriptstyle\pm.03$ 0.55±.08plus-or-minus0.55.080.$5$08 $0.82\scriptstyle\pm.04$ $0.56\scriptstyle\pm.08$ $0.79\scriptstyle\pm.05$
Mb+Ci $0.47\scriptstyle\pm.12$ $0.68\scriptstyle\pm.08$ $0.58\scriptstyle\pm.09$ $0.71\scriptstyle\pm.06$ 0.77±.05plus-or-minus0.77.050.$7$05 0.88±.03plus-or-minus0.88.030.$8$03 $0.79\scriptstyle\pm.04$ 0.88±.02plus-or-minus0.88.020.$8$02
[…]
Cdn (no $\mathcal{L}_{G}$) $0.65\scriptstyle\pm.07$ $0.86\scriptstyle\pm.06$ $0.72\scriptstyle\pm.09$ $0.84\scriptstyle\pm.07$ $0.81\scriptstyle\pm.05$ $0.95\scriptstyle\pm.02$ $0.86\scriptstyle\pm.03$ $0.94\scriptstyle\pm.01$
Soft Ut-Igsp $10$01 $0.69\scriptstyle\pm.03$ $0.19\scriptstyle\pm.01$ $0.56\scriptstyle\pm.03$ $0.18\scriptstyle\pm.02$ 0.77±.02plus-or-minus0.77.020.$7$02 0.22±.01plus-or-minus0.22.010.$2$01 $0.60\scriptstyle\pm.01$
Dcdi-G $0.17\scriptstyle\pm.03$ $50$05 $0.23\scriptstyle\pm.01$ 0.44±.03plus-or-minus0.44.030.$4$03 $0.14\scriptstyle\pm.03$ $0.48\scriptstyle\pm.03$ $0.30\scriptstyle\pm.04$ $0.54\scriptstyle\pm.04$
Dcdi-Dsf $0.17\scriptstyle\pm.04$ $0.46\scriptstyle\pm.04$ $0.24\scriptstyle\pm.02$ $0.47\scriptstyle\pm.03$ $0.16\scriptstyle\pm.02$ $50$05 $0.26\scriptstyle\pm.02$ $0.48\scriptstyle\pm.02$
Bacadi-E $0.25\scriptstyle\pm.07$ $0.65\scriptstyle\pm.03$ $0.40\scriptstyle\pm.13$ $0.68\scriptstyle\pm.08$ $0.43\scriptstyle\pm.22$ $0.81\scriptstyle\pm.09$ $0.49\scriptstyle\pm.15$ 0.77±.05plus-or-minus0.77.050.$7$05
Bacadi-M 0.11±.01plus-or-minus0.11.010.$1$01 $0.61\scriptstyle\pm.02$ $0.26\scriptstyle\pm.07$ $0.64\scriptstyle\pm.06$ $0.25\scriptstyle\pm.10$ $0.80\scriptstyle\pm.09$ $0.38\scriptstyle\pm.10$ $0.76\scriptstyle\pm.06$
Dci $0.48\scriptstyle\pm.02$ $0.78\scriptstyle\pm.02$ $0.53\scriptstyle\pm.06$ 0.77±.04plus-or-minus0.77.040.$7$04 $0.62\scriptstyle\pm.09$ $0.89\scriptstyle\pm.04$ 0.55±.04plus-or-minus0.55.040.$5$04 $0.80\scriptstyle\pm.04$
Mb+Ci $0.46\scriptstyle\pm.10$ $0.65\scriptstyle\pm.06$ $0.49\scriptstyle\pm.04$ $0.64\scriptstyle\pm.03$ $\textbf{0.97}\scriptstyle\pm.02$ $0.98\scriptstyle\pm.01$ $0.92\scriptstyle\pm.05$ $0.94\scriptstyle\pm.03$
@@
Baselines  We compare against discrete and continuous causal discovery algorithms for unknown interventions. Ut-Igsp (Squires et al., 2020) infers causal graphs and unknown targets by greedily selecting the permutation of variables that minimizes their proposed score function. Dcdi (Brouillard et al., 2020) and Bacadi (Hägele et al., 2023) are continuous causal discovery algorithms that fit generative models to the data, where the causal graph and intervention targets are model parameters. Dcdi’s -G and -Dsf suffixes correspond to Gaussian and deep sigmoidal flow parametrizations of the likelihood. Bacadi’s -E and -M suffixes indicate empirical (standard) and mixture (bootstrap) variants.
Mb+Ci is inspired by Huang et al. (2020). We introduce a domain index (indicator of environment), compute its Markov boundary using graph lasso, and use conditional mutual information to identify true children of the domain index (intervention targets). Dci with stability selection (Belyaeva et al., 2021) predicts edge-level differences between two causal graphs. We take each node’s proportion of changed edges as its likelihood of being an intervention target. While Dci was also motivated by biological applications, it only scales to around a hundred variables at most, so we evaluate it alongside other causal discovery methods here.
Evaluation  In the synthetic case, we are not constrained by biological redundancy or incomplete predictions, so we report standard classification metrics: mean average precision (mAP) and area under the ROC curve (AUC). Both metrics are computed independently for each variable and averaged over all regimes of the same number of targets. Their values range from 0 to 1 (perfect). The mAP random baseline depends on the positive rate, while the AUC random baseline is 0.5 (per edge).
Results  On synthetic data, Cdn achieves high performance across intervention types and data-generating mechanisms (Table 3), while running in seconds (Table 11). As an ablation study, we investigated removing the graph loss and the pretrained weights, with no change to the graph-centric architecture. We found that supervising based on the graph is almost always helpful (“no $\mathcal{L}_{G}$” row), and pretraining leads to more stable training dynamics (Figure 10).
@@
Dci and Mb+Ci are the best baselines, which is encouraging, as they are designed with the intuition that detecting differences is easier than reproducing the entire causal graph. Surprisingly, most other baselines perform poorly at recovering intervention targets. In the case of Dcdi and BaCaDI, this may be because it is hard to select a single sparsity threshold for varying sizes of intervention sets, and to balance the sparsity regularizer with the generative modeling objective.
Impact statement
Our work proposes a machine learning method to improve understanding of large-scale perturbation experiments, in the context of molecular biology. Towards broader impact, this work may facilitate better understanding of drug mechanisms; inform protocol design for cellular reprogramming; or identify causal targets of disease. While this work has numerous applications to the life sciences, it does not directly concern the design of molecules or other potentially harmful agents.
@@ Acknowledgements
Appendix A Model architecture
A.1 Attention layers
To balance expressivity and efficiency, our model is designed around series of “axial” attention layers, similar to Ho et al. (2020); Rao et al. (2021); Wu et al. (2024). The causal structure learner closely resembles the architecture of Sea (Wu et al., 2024), while the differential network is slightly different (since the input space is different). Briefly, the input to the causal structure learner is a matrix of shape $N\times N\times T\times d$. However, the (dense) global statistic does not span the $T$ dimension, and the local graph estimates are generally sparse. Thus, the causal graph learner separately attends over the $N\times N\times d$ and $K\times T\times d$ components ($K$ is the number of unique edges in sparse format), with “messages” to align the representations between each attention layer. In contrast, the differential network only operates over the dense $N\times(N+1)\times d$ (or $2d$) representation. Thus, it omits the message passing layers and directly stacks dense attention layers.
A.2 Edge embeddings
Since attention itself is invariant to input ordering, Transformer-style architectures use absolute (Devlin et al., 2019) or relative (Su et al., 2024) positional embeddings to distinguish input elements. Images and sequences have a natural ordering among their elements. This structure inherently induces distances between positions, which provides signal for learning models. In contrast, models over graphs should be equivariant to the labeling of nodes. That is, if $1\to 2$ is an edge, while $2$ and $3$ are not connected, then re-labeling the nodes $(1,2,3)\mapsto(2,3,1)$ should result in edge $2\to 3$ and no edge between $3$ and $1$.
If we assign an arbitrary ordering to each graphx, the Transformer will be able to distinguish between nodes. However, there are several practical issues. Since synthetic data are generated in the topological ordering of the nodes, always assigning the root nodes to low positions may lead to unintended information leakage. Furthermore, if the model is trained on graphs of up to $N=100$, but tested on larger graphs, the positional embeddings associated with $N>100$ will be entirely random and out of distribution. Following Wu et al. (2024), we randomly sample a permutation of the maximum graph size to each graph. Each edge embedding is the concatenation of its constituent nodes’ embeddings, combined via a feed-forward network:
$\text{Embed}(i,j)=\text{FFN}([\text{Embed}(i),\text{Embed}(j)])$ (9)
where $i,j$ are arbitrarily assigned, but consistent within the graph. This embedding is added to the hidden representation $h_{i,j}$ before the first layer of attention.
In the differential network, the additional “column” associated with node-level statistics is not subject to any embedding (which implicitly differentiates it from the remaining inputs), while the “row” identifies the node these statistics were computed for.
Appendix B Implementation details
B.1 Cdn training
We trained until convergence (no improvement for 50 epochs on synthetic; 10 epochs on Perturb-seq) and selected the best model based on validation mAP (Figure 4). The validation distribution was identical to training and did not contain test causal mechanisms. During training, we used 15 CPU workers (primarily for local graph estimates) and 1 A6000 GPU. Cdn took around 6 hours to train on synthetic data, and 1 hour to finetune on real data. Unless otherwise noted, the causal structure learner was initialized from the Sea inverse covariance, FCI pretrained weights (Wu et al., 2024).
Figure 4: Hyperparameter search: number of differential network layers. We selected 3 layers based on validation metrics.
B.2 Baselines
We used the latest releases of all baselines.
Gears is a graph neural network that predicts the effects of unseen genetic perturbations, based on the Gene Ontology knowledge graph (Ashburner et al., 2000). We trained Gears on each cell line separately, and predicted the effects of perturbing every expressed gene. Then we ranked candidates based on cosine similarity to the interventional data. Figure 9 depicts the low, but non-trivial proportion of differentially expressed genes for which Gears was unable to make a prediction, due to lack of node coverage in their processed gene ontology graph (Ashburner et al., 2000). These genes were not considered in the rankings or evaluations for Gears.
GenePt utilized the v2 March 2024 update, which used newer models and additional protein data, compared to their initial paper (GenePT_gene_protein_embedding_model_3_text). We concatenated each perturbation’s gene embedding to the log-fold change of the top 5000 highly-variable genes (Wolf et al., 2018) and trained a logistic regression model on each cell line to predict true vs. decoy perturbations. Candidates were ranked based on predicted probability. Of the 2,842 unique genetic perturbations, only 2,819 ($99.1\%$) mapped to GenePt embeddings. The remainder used the mean gene embedding (within the dataset) as the language-based embedding.
Ut-Igsp used the official implementation from commit 7349396, since the repository is not back-compatible.
Bacadi is evaluated using the fully-connected implementation, since it performed better than the linear version. However, since the linear version is significantly faster, we include its runtime in Table 11.
Mb+Ci was implemented using scikit-learn (Pedregosa et al., 2011).
Appendix C Theoretical motivation
While this paper focuses on biological applications, we show that the differential network is well-specified as a class of models, i.e. a correct solution is contained within the space of possible parametrizations.
@@ Appendix C Theoretical motivation
In this analysis, we assume that the “causal” representations $h$ contains enough information both to retain global statistics and recover the true graph $E$. The former is reasonable due to high model capacity, and the latter is based on high empirical performance in graph reconstruction (Wu et al., 2024). We emphasize that the latter is an empirical judgment, which may not hold on all datasets in practice.
Attention-based architecture
Our differential network is implemented using an attention layer, which is composed of two self-attention layers (one along each axis of the adjacency matrix) and a feed-forward network. For simplicity, we follow prior work (Yun et al., 2019) and ignore layer normalization and dropout.
Our inputs $h\in\mathbb{R}^{2d\times N\times N}$ are $2d$-dimension features, which represent a pair of $N\times N$ causal graphs. We use $h_{\cdot,j}$ to denote a length $N$ row for a fixed column $j$, and $h_{i,\cdot}$ to denote a length $N$ column for a fixed row $i$. The attention layer implements:
@@
Here, the attention layer should implement $h-h^{\prime}$, so that when we collapse over the incoming edges, the output is non-zero only at the edge differences. Suppose the first dimension of the $2d$ feature stores $E$, and the second dimension stores $E^{\prime}$. The row self-attention implements the identity (in the first two dimensions). Then we can set $W_{K,Q}$ to zero, $W_{V}$ to the identity, and $W_{O}$ to
@@
where $c_{i}$ are sampled at random per synthetic dataset. This choice is inspired by the fact that biological perturbation effects are measured in fold-change. Here, the adjacency matrices are the same, but global statistics differ. In particular, we focus on two statistics: the correlation matrix $R$ and the covariance matrix $\Sigma$. Note that while we do not explicitly provide the covariance matrix as input, we do provide marginal variances, from which the same information can be derived.
@@
Appendix D Datasets
D.1 Biological datasets
We converted all single cell datasets to log-normalized, transcripts per 10,000 UMIs. Perturb-seq genes were mapped to standard identifiers and filtered by the authors. Sci-Plex dataset variables represented genes that appeared in at least 5,000 cells (threshold chosen to achieve a similar number of genes). We performed differential expression analysis via the scanpy package (Wolf et al., 2018), using the Wilcoxon signed-rank test (Wilcoxon, 1945) with Benjamini-Hochberg p-value correction and a threshold of adjusted p-value $<0.05$. Table 4 reports statistics of the raw, unprocessed datasets.
For Perturb-seq datasets, we kept perturbations with $>10$ differentially-expressed genes (DEGs), and clustered them using k-means with $k=200$, chosen heuristically based on log-fold change heatmaps (Figure 5). These clusters were used to inform data splits for seen cell lines, where the largest cluster was allocated to the training set, and all remaining clusters were split equally among train and test. The largest cluster(s) appear to contain perturbations with smaller effects. For Sci-Plex datasets, only 6 drug perturbations across 2 cell lines resulted in differential expression of their known protein targets (Supplementary Table 8 from McFaline-Figueroa et al. (2024)). Therefore, we used these exclusively as test sets. Table 5 reports statistics of the final, processed datasets. Figure 8 plots the full distribution of number of cells and DEGs per perturbation.
Figures 6 and 7 depict the differences in gene co-expression matrices between control and perturbed cells. Genetic perturbations tend to have much clearer phenotypes compared to chemical perturbations, whose effects are more diffuse.
The full names of the Sci-Plex drugs and their targets are as follows. These targets were collated by the authors (McFaline-Figueroa et al., 2024), and are each are well-documented in the literature.
  • Infigratinib (“infig”) binds target FGFR1 with nano-molar affinities (Knox et al., 2024).
  • Nintedanib (“nint”) and target FGFR1 have been probed through structural studies (Hilberg et al., 2008).
  • Palbociclib (“palb”) has been co-crystallized (PDB 5L2I) with its target CDK6.
  • Doxorubicin (“doxo”) and target TOP2A are associated in a number of works (Knox et al., 2024).
  • Volasertib (“vola”) has been co-crystallized (PDB 3FC2) with PLK1 and a number of works confirm specificity for PLK2 (Chen et al., 2021).
It is important to note that these drugs are also associated with known off-targets. However, we did not observe differential expression of any off-target, adjusted $p\approx 1$, perhaps because the impact on their mRNA expression is negligible. Protein activity would be more suitable for quantifying the impact of a drug on a cell. However, this cannot be measured at scale. In fact, even protein abundance is difficult to measure at single-cell resolution, as it relies on mass spectrometry to differentiate proteins (Newman et al., 2006), while mRNA can be exhaustively enumerated and sequenced (based on the known human genome). Therefore, this chemical perturbation study is primarily a proof of concept, to probe whether targets can be inferred from mRNA expression, and if so, to what extent.